In this assignment, you are asked to write a matlab program to solve the following open ended problem, and submit the program and written summary including methods employed, algorithm implementation, final results and recommendations on further improvement.

Background

A drug administered to a patient produces a concentration in the blood stream given by

                                               C(t)=Atexp(-t/3 )

Where c is in milligrams per millimeter, t is in hours and A is injected units. For a given drug the maximum safe concentration (in mg/ml) is denoted as cmax, and the critical concentration(in mg/ml), denoted as ccritical, is the concentration below which an additional amount of drug needs to be administered to the patient.

a.      Determine how the injected unit A affects the maximum safe concentration and when the maximum safe concentration occurs.

b.      Determine how the ratio between ccritical and cmax affects the duration between the first and the second injections.

Suggestions



1
Expert's answer
2021-11-01T16:23:19-0400

a.

"C'(t)=Ae^{-t\/3}-Ate^{-t\/3}\/3=0"

"1-t\/3=0"

"t=3" hours

"C_{max}=C(3)=3A\/e=1.1A"


b.

"C_{critical}=te^{-t\/3}"

"C_{max}=2\cdot 1.1=2.2"

"C_{critical}\/C_{max}=te^{-t\/3}\/2.2"